### v1

# LocationReasoner: Evaluating LLMs on Real-World Site Selection Reasoning

Abstract
Recent advances in large language models (LLMs), particularly those enhanced through reinforced post-training, have demonstrated impressive reasoning capabilities, as exemplified by models such as OpenAI o1 and DeepSeek-R1. However, these capabilities are predominantly benchmarked on domains like mathematical problem solving and code generation—leaving open the question of whether such reasoning skills generalize to complex, real-world scenarios. In this paper, we introduce LocationReasoner, a benchmark designed to evaluate LLMs’ reasoning abilities in the context of real-world site selection, where models must identify feasible locations by reasoning over diverse and complicated spatial, environmental, and logistical constraints. The benchmark comprises over 300 carefully crafted queries of varying difficulty levels, supported by a sandbox environment with in-house tools for constraint-based location search. Extensive evaluations reveal that state-of-the-art reasoning models offer limited improvement over their non-reasoning predecessors in real-world contexts, with even the latest OpenAI o4 model failing on 30% of site selection tasks. Moreover, agentic strategies such as ReAct and Reflexion often suffer from over-reasoning, leading to worse outcomes than direct code-generation prompting. With key limitations of LLMs in holistic and non-linear reasoning highlighted, we release LocationReasoner to foster the development of LLMs and agents capable of robust, grounded reasoning in real-world decision-making tasks. Codes and data for our benchmark are available at https://github.com/miho-koda/LocationReasoner.
1 Introduction
Large language models (LLMs) have achieved tremendous success across a broad spectrum of tasks, ranging from human-like dialogue achiam2023gpt ; team2023gemini ; liu2024deepseek ; bai2023qwen ; glm2024chatglm ; touvron2023llama to agent-based simulations park2023generative ; shanahan2023role ; wang2024survey –progress largely driven by scaling data and compute sutton2019bitter ; wei2022emergent . Recent advances extend the notion of scaling beyond training, introducing inference-time scaling techniques such as chain-of-thought (CoT) prompting wei2022chain and agentic strategies like ReAct yao2023react , which enable deeper reasoning through intermediate steps and tool use. In particular, large-scale reinforcement learning has been applied during post-training to generate high-quality reasoning traces (e.g., CoT), giving rise to a new class of reasoning models such as OpenAI o1 jaech2024openai and DeepSeek-R1 guo2025deepseek . These reasoning capabilities are critical for enabling LLMs to tackle complex, real-world tasks that often require decomposing problems, exploring alternative strategies, and interacting with external tools.
Benchmarks play a crucial role in driving the progress of LLMs patterson2012better , and the rapid advancement of these models has created a growing demand for more challenging and informative evaluation tasks. While general language understanding benchmarks are widespread chiang2024chatbot ; wang2024mmlu ; hendrycks2021mmlu ; rein2024gpqa ; huang2023c , assessing reasoning capabilities remains considerably more difficult—primarily due to the challenge of verifying the correctness of LLMs’ reasoning outputs. As a result, current reasoning models are predominantly evaluated on domains with easily verifiable solutions, where verifying a solution is much simpler than generating one. This includes domains such as mathematical problem solving and programming, exemplified by benchmarks like AIME aime2024 , MATH500 math500 , and CodeForces quan2025codeelo . In contrast, few benchmarks have tackled real-world reasoning scenarios, which often involve more complicated constraints than those found in math or coding. For instance, the TravelPlanner benchmark xie2024travelplanner evaluates LLM agents on practical planning tasks but does not include recent reasoning models and lacks automated verification mechanisms, relying heavily on manual annotation. Consequently, it remains unclear whether current reasoning LLMs and agents can effectively generalize to real-world reasoning tasks. This gap highlights the urgent need for a new benchmark that targets real-world reasoning while remaining scalable and annotation-free.
In this work, we introduce LocationReasoner, a benchmark designed to evaluate the real-world reasoning capabilities of LLMs through the task of site selection. Site selection is a common decision-making process in practical business settings, where the objective is to identify feasible locations based on a diverse set of criteria. This task presents a significant reasoning challenge, as it requires LLMs to perform structured location search and filtering over queries that involve multiple constraints with complex logical dependencies. For instance, selecting a site for a new restaurant may demand reasoning over temporal consumption constraints and spatial transportation constraints, with various conditions often combined through Boolean logic. To comprehensively assess LLMs’ ability to reason under such practical and multifaceted conditions, we construct the LocationReasoner benchmark, which includes over 300 site selection queries spanning a wide range of difficulty levels. The benchmark is supported by a sandbox environment equipped with offline datasets and in-house tools for constraint-based location search.
LocationReasoner differs fundamentally from existing benchmarks in two key aspects. First, it focuses on practical reasoning, where LLMs must decompose complex constraints into smaller executable steps and generate codes using available tools to solve each query. Second, the benchmark features deterministic and automatically verifiable queries, enabling fully scalable evaluation without human annotation, and allowing seamless extension to much larger benchmark suites.
We summarize the key contributions of this work as follows:
  • We introduce LocationReasoner, a real-world reasoning benchmark that grounds LLMs in practical site selection tasks. The benchmark includes a suite of curated datasets, toolsets, and a sandbox environment to enable out-of-the-box evaluation of LLMs in complex constraint-based reasoning.
  • We conduct a comprehensive evaluation of four major LLM families–OpenAI jaech2024openai , Gemini team2023gemini , Claude anthropic2024claude35sonnet , and DeepSeekguo2025deepseek –covering both general-purpose and reasoning-augmented models, and two representative agentic workflows, ReAct yao2023react and Reflexion shinn2023reflexion .
  • Our benchmarking results reveal that current LLMs struggle to effectively handle real-world reasoning challenges. For instance, the latest OpenAI o4 model achieves only 69.99% success on site selection queries. Furthermore, reasoning-augmented models and agentic strategies provide limited gains compared to direct prompting of general models for code generation. In-depth analysis uncovers key limitations in holistic and non-linear reasoning, pointing to critical bottlenecks that need to be addressed to improve LLM performance on practical, real-world tasks.
2 LocationReasoner
Figure 1: Query generation and execution pipeline. Queries are automatically generated using predefined constraints, ranges, thresholds, and logic composition—either through rule-based or LLM-based methods. Each query is executed via a deterministic code path using interdependent in-house tools or through an LLM agent. The results are then compared to evaluate whether the agent produces the correct filtered zones.
We center the LocationReasoner benchmark around practical site selection, a problem that is easy to understand and intuitive, but also requires multi-step reasoning across spacial, economic, and infrastructural constraints. Built on real data from Boston, MA and a fixed set of in-house tools, the benchmark offers a controlled and interpretable environment to test LLMs’ ability to reason and plan under real-world constraints.
As illustrated in Figure 1, we construct the entire query generation and execution pipeline such that it can be fully automated to support scalable evaluation. By programmatically defining constraints, threshold ranges, and logical compositions, queries can be generated in bulk through either rule-based generation or LLM-based generation. The rule-based approach is rigid and syntactically precise as predefined constraints are directly encoded into executable templates. It ensures that all conditions are logically well-formed and interpretable by deterministic code. LLM-based generation introduces linguistic diversity. The system feeds the same set of structured constraints to a language model and asks it to produce free-form queries in natural language. These queries are then routed through two execution pathways: a deterministic code-based system that applies filters using in-house tools, and an LLM agent system to be evaluated that interprets and solves the same queries. The system logs and compares the final site selections from both paths, enabling large-scale benchmarking without manual annotation or intervention.
2.1 Environment Setting
We construct a sandbox environment to ensure a stable and consistent evaluation of all LLM agents, where datasets are fixed, no external API calls are made, and all tool functionalities remain constant.
Datasets. LocationReasoner utilizes data from Boston, MA for site selection. The benchmark database is constructed by integrating multiple real-world datasets to provide a rich and realistic foundation for constraint-based planning. Specifically, we construct the database by merging data from the SafeGraph dataset [^1] https://www.safegraph.com/, which includes detailed information on Points of Interest (POIs), parking facilities, and consumer spending patterns from 2019 to 2025. To enrich the spatial and demographic context, we incorporate population data from Google Places API [^2] https://developers.google.com/maps/documentation/places/web-service/overview and transportation network data from OpenStreetMap [^3] https://www.openstreetmap.org/. This integrated dataset enables agents to reason effectively about accessibility, mobility, and demand when evaluating site suitability.
Tools. To support structured reasoning, the sandbox contains a set of in-house tools that LLM agents can call during problem solving. These tools expose curated functionalities such as filtering zones, analyzing parking availability, retrieving spending patterns across time, and evaluating transportation accessibility. We provide a detailed description of each tool and its inputs/outputs in Table 1.
Table 1: In-house tools description.
Category Tool Name Description
Loader Tools get_poi_spend_dataset Loads POI metadata and consumer spending patterns.
get_parking_dataset Loads parking facility data including location and area.
Zone Tools create_zone Creates a zone DataFrame of POIs grouped by zone.
assign_parking_zones Assigns each parking lot to a zone.
get_zone_center Returns the geographic center of a zone.
get_neighbor_zones Finds the closest zones to a given zone by distance.
Analysis Tools get_spendparam_years Computes aggregated spend over selected years.
get_num_parking Counts parking lots per zone.
get_largest_parking_lot_area Returns the area of the largest parking lot per zone.
get_largest_parking_capacity Returns the maximum parking capacity per zone.
get_distance_km Calculates haversine distance between two points.
Filter Tools filter_df_based_on_zone Filters a DataFrame to only include specified zones.
filter_pois_by_top_category Filters POIs by top-level category.
filter_pois_by_sub_category Filters POIs by sub-category.
get_transport_pois_in_zone Finds transport POIs by type and groups them by zone.
Population Tool get_population Retrieves population per zone using Google Places API.
2.2 Query Design
We design queries to systematically evaluate an LLM agent’s ability to reason over various constraints using in-house tools. Queries are categorized into three difficulty levels, reflecting an increasing level of complexity. Difficulty is determined by the number of constraints, diversity of tools involved, the depth of reasoning required, the linguistic complexity of the query, and the necessity to synthesize and compare heterogeneous data sources across spatial and temporal dimensions. The difficulty level enables controlled benchmarking across different kinds of reasoning scenarios, ranging from straightforward constraint filtering to complex, multi-faceted decision-making. The query variants mimic the types of natural, often nuanced questions that real users might ask when selecting a site, testing the LLM’s ability to precisely extract and interpret user requests. We provide detailed descriptions and examples of different difficulty levels as follows:
  • Easy (114 queries). Easy queries involve direct calls to a single in-house tool with a simple numerical threshold or binary filter. These queries require minimal coordination between tools.
    Example: “I want to build a retro roller rink. But the zone needs at least 3 parking lots. Where should I look?”
  • Medium (102 queries). Queries at this level introduce logical dependencies between constraints and often require coordination across multiple tools. These queries combine 2 to 3 simple constraints using Boolean operators such as AND or OR, chaining together multiple easy-level conditions. Additionally, medium queries demand mathematical reasoning and often involve relationships across structural features (e.g., neighboring zones), socio-economic indicators (e.g., spending patterns), or temporal variation (e.g., comparing data from different years).
    Example: "Find zones where at least 35% of total raw total spend in 2020 comes from the top category {Restaurants and Other Eating Places}."
  • Hard (102 queries). Hard queries consist of typically 3-6 composite constraints derived from simple or medium-level logic, combined using AND, OR, and NOT operators. These queries demand nuanced, multi-tool reasoning, trade-off analysis, and often involve exclusionary conditions. They closely resemble real-world decision-making scenarios in site selection, where users express both strong preferences and explicit disqualifiers, requiring the agent to balance multiple objectives while avoiding infeasible options.
    Example: "Launch a creative co-working café — needs at least 26 POIs total in that area and also strong local spending with 50%+ from sub-category of {Full-Service Restaurants} in 2022, but I’m not interested if the number of POIs in that same category dominates the area by 30%."
2.3 Reasoning Approaches
Our goal is to assess whether LLMs can reason over multi-step planning constraints, invoke appropriate in-house tools, and return correct site selection outputs. We explore three distinct strategies:
  • Direct Code Generation. We provide LLMs with relevant context of available in-house tools and structured data, prompting it to directly generate executable Python code that fulfills the query.
  • ReAct (Reasoning + Acting). The ReAct framework yao2023react is adopted to guide the agent through an iterative loop of Thought-Action-Observation (TAO). In the Thought step, the agent reflects on its current state and reasons about what to do next. In the Action step, it either calls an in-house tool or runs custom Python code. Executing custom Python code is particularly useful for performing transitional logic such as filtering results, combining intermediate outputs, or applying custom calculations. The resulting output is captured in the Observation step and stored as part of the agent’s evolving internal state. These observations are continuously fed back to the model, enabling it to reason with an evolving memory of prior tool outputs and decisions. The model is also allowed to store and reference intermediate results.
  • ReAct + Reflexion. Reflexion shinn2023reflexion extends the ReAct framework by allowing the model to reflect on failed attempts such as code errors or reaching the maximum step limit without producing a result. In these cases, the model receives feedback from its prior reasoning trace and uses it to revise its approach. This added layer of self-correction helps the model identify mistakes, adjust its logic, and improve tool usage in subsequent retries, making it especially useful for handling complex, multi-constraint queries.
2.4 Automated Verification
To ensure that each test case has a clearly defined and feasible solution, we develop a set of deterministic, rule-based functions that serve as the ground truth logic for evaluating site selection. These logic scripts are constructed using the same in-house tools available to the LLM agent and are designed to satisfy all constraints specified in each query. For each query, the script returns a set of valid zones that fully meet the defined conditions. During evaluation, the output of the LLM agent is compared directly against these ground truth results to assess correctness. This approach ensures consistent, reproducible evaluation across all queries, and is automated without any human annotation. We evaluate the agent’s ability to generate complete outputs, satisfy user-defined constraints, and accurately identify valid zones with the following complementary metrics:
  • Delivery Rate. Delivery Rate measures the proportion of queries for which the LLM agent successfully returns selected zones. A process is considered completed if the agent successfully produces an output without execution failure. For agents evaluated via direct code generation, the generated code must run without syntax or runtime errors. For agents evaluated under the ReAct or Reflexion framework, the agent must avoid generating incorrect or unexecutable code, prevent infinite reasoning loops, and successfully complete the task within a strict limit of 30 reasoning steps. This metric captures basic functional reliability, as failed plans cannot be evaluated for correctness.
  • Perfect Pass Rate. The Perfect Pass Rate measures the percentage of outputs in which the set of zones returned by the agent exactly matches the ground truth derived from our objective logic. A plan is considered perfect only if it satisfies all constraints and produces no extraneous or missing zones. This strict metric ensures full compliance with the query’s requirements and serves as a high-confidence indicator of planning accuracy.
  • Precision, Recall, and F1 Score. In addition to binary success metrics, we evaluate zone-level selection accuracy using standard classification metrics. Precision measures the fraction of zones selected by the agent that are actually valid according to the ground truth. Recall measures the fraction of valid zones from the ground truth that the agent successfully identified. F1 Score is the harmonic mean of precision and recall, offering a balanced view of over- and under-selection.
3 Experiments
For the direct code generation setting, we test a range of models spanning multiple providers, including DeepSeek (V3, R1), OpenAI (GPT-4o, o4-mini), Gemini (1.5, 2.5), Claude (3, 3.5). Each model receives the same prompt structure, tool descriptions, and structured input data, and is tasked with generating executable Python code that satisfies the constraints specified in each query. Notably, we cover both general and reasoning models of each LLM family, to assess whether such advertised reasoning capabilities lead to improved performance in real-world structured, constraint-driven reasoning tasks. For the ReAct and Reflexion frameworks, we focus our evaluation on OpenAI’s GPT-4o model. This setup allows us to examine the benefits of iterative reasoning, tool chaining, and self-correction over single-shot code generation approaches.
3.1 Main Results
Table 2: Performance metrics on LocationReasoner by model and difficulty level.
Model Difficulty Delivery Rate Perfect Pass Precision Recall F1 Score
Code Generation
Deepseek V3 simple 92.98% 68.42% 0.73 0.69 0.67
medium 79.17% 60.42% 0.55 0.52 0.51
hard 77.19% 42.11% 0.24 0.20 0.18
overall 83.33% 57.00% 0.51 0.47 0.46
Deepseek R1 simple 92.59% 75.00% 0.73 0.71 0.69
medium 76.04% 61.46% 0.55 0.54 0.51
hard 71.57% 36.27% 0.23 0.25 0.22
overall 80.39% 58.92% 0.51 0.51 0.48
simple 89.81% 71.30% 0.66 0.61 0.60
medium 85.42% 55.21% 0.57 0.46 0.45
hard 88.24% 37.25% 0.26 0.17 0.15
OpenAI 4o overall 87.91% 55.00% 0.50 0.42 0.40
simple 91.82% 81.82% 0.74 0.74 0.72
medium 89.58% 73.96% 0.67 0.64 0.61
hard 88.24% 53.92% 0.42 0.40 0.38
OpenAI o4-mini overall 89.94% 69.99% 0.61 0.59 0.57
Gemini 1.5 simple 46.30% 29.63% 0.32 0.30 0.30
medium 51.04% 32.29% 0.27 0.25 0.23
hard 46.08% 19.61% 0.13 0.08 0.09
overall 47.71% 27.00% 0.24 0.21 0.20
Gemini 2.5 simple 92.59% 77.78% 0.65 0.65 0.64
medium 80.21% 63.54% 0.52 0.50 0.48
hard 83.33% 45.10% 0.33 0.35 0.32
overall 85.62% 62.00% 0.50 0.51 0.48
simple 66.67% 35.19% 0.39 0.34 0.31
medium 47.92% 21.88% 0.14 0.08 0.07
hard 30.39% 9.80% 0.03 0.01 0.01
Claude 3 Haiku overall 48.69% 23.00% 0.19 0.15 0.13
simple 67.59% 52.78% 0.51 0.50 0.48
medium 65.62% 54.17% 0.43 0.40 0.39
hard 46.08% 26.47% 0.22 0.21 0.18
Claude 3.5 Haiku overall 59.80% 44.00% 0.39 0.37 0.35
Agentic Workflow
ReAct simple 72.22% 39.81% 0.44 0.35 0.36
medium 92.71% 26.04% 0.37 0.17 0.18
hard 85.29% 16.67% 0.25 0.07 0.09
overall 83.01% 28.00% 0.35 0.20 0.21
ReAct + Reflexion simple 99.07% 32.41% 0.54 0.27 0.29
medium 94.79% 37.50% 0.52 0.29 0.31
hard 97.06% 29.41% 0.31 0.13 0.14
overall 97.06% 33.00% 0.46 0.23 0.24
Limited overall performance. The overall model performance on the LocationReasoner benchmark remains limited, as most models fail to generate complete and accurate outputs across varying difficulty levels. The average perfect pass rate across 8 different LLMs is only 49.61%. Even the best-performing model, OpenAI o4-mini, achieves only a 69.99% perfect pass rate and an overall F1 score of 0.57. This reflects a significant gap between current LLM capabilities and the demands of real-world complex reasoning tasks.
Difficulty sensitivity. Across all models, performance consistently degrades as task difficulty increases. All key metrics decline from simple to medium to hard queries. Specifically, the average perfect pass rate on hard queries is only 33.82%, which is much lower than 61.49% of easy queries. This trend highlights that while some LLMs may be capable of handling basic filtering or single-constraint tasks, they struggle to generalize to multi-step reasoning and more compositional decision-making challenges.
Reasoning LLMs offer significant improvements. Despite the DeepSeek family, we observe significant progress in all other three families made by their reasoning models compared with their baseline counterparts. For example, Gemini 2.5, advertised as a reasoning-optimized model, achieves an overall F1 score of 0.48 and a perfect pass rate of 62%. Its predecessor, Gemini 1.5, reaches only a 27% perfect pass rate and an F1 score of 0.21. However, the gains remain far from transformative. Gemini 2.5 still fails over 35% of simple queries and more than half of medium ones. This underscores that even when reasoning-optimized models do improve over prior versions, their performance still falls short of the reliability required for real-world reasoning tasks.
Agentic strategies do not guarantee better results. Agentic strategies such as ReAct and Reflexion do not outperform direct code generation. Despite being designed to simulate step-by-step human reasoning, ReAct + Reflexion achieves only a 33% perfect pass rate and an overall F1 score of 0.24, which is substantially lower than the best-performing code generation LLMs. This suggests that agentic stratety alone is insufficient without more reliable reasoning abilities of the base model. We provide detailed explanations on the inferior performance of agentic strategies in Appendix A.
4 Error Analysis
Agents consistently struggle with five primary failure types, each corresponding to the categories shown in Figure 2. These errors often overlap, but their root causes reflect distinct reasoning weaknesses that hinder constraint satisfaction and plan validity.
Figure 2: Breakdown of failure types for OpenAI and DeepSeek models
  • Logic Error. Logic errors encompass broken reasoning chains, incorrect constraint handling, and improperly ordered filtering logic. This includes sequential filtering mistakes where each constraint is applied too early, altering the input for future conditions. In ReAct, logic errors often emerge as inconsistent multi-step chains, where prior outputs are forgotten or overwritten. In code generation, they tend to surface as incorrect condition checks or missing final evaluations.
  • Edge Case Error. LLMs frequently misinterpret constraints that require careful logic around null values or boundary conditions. A common example involves requests like “fewer than 3 competitors,” where agents exclude zones with zero competitors–despite these being valid. These failures stem from inflexible logic structures and an inability to reason inclusively.
  • Tool Error. This category captures failures where the model misuses, miscalls, or entirely ignores in-house tools. Examples include passing incorrect argument formats, selecting a tool misaligned with the constraint type, or chaining incompatible tool outputs. Code generation sometimes bypass the toolset and attempt to write their own versions of utility functions, which introduces further bugs and inconsistencies. These issues reveal brittle tool affordance understanding and poor mapping between language and tool usage.
  • Prompt Error. Prompt-related failures involve misinterpreting user intent. In many cases, they misclassify the type of reasoning needed (e.g., assuming a ranking is needed when only filtering is requested), leading to logic or tool selection failures. These mistakes reflect weak grounding in language understanding and sensitivity to phrasing variations.
  • Code Error. Code-specific issues include syntax errors, missing return statements, incorrect indentation, and unsafe operations such as division by zero. These errors prevent valid plans from executing altogether and reduces the agent’s delivery rate.
Table 3: Correction rates (%) for reasoning vs. non-reasoning model pairs across difficulty levels.
Difficulty OpenAI o4-mini vs 4o Claude 3.5 vs 3 Deepseek R1 vs V3 Gemini 2.5 vs 1.5 ReAct Reflexion vs None
simple 18.89 28.26 10.10 21.74 6.67
medium 23.68 30.30 2.82 20.00 18.82
hard 17.95 16.67 4.08 21.62 16.87
We evaluate whether reasoning models can correct the errors made by their nonreasoning counterparts. Table 3 shows the results where Correction Rate is defined as the proportion of test cases in which the reasoning model produces a correct output while the non-reasoning model fails. Our results indicate that reasoning models are particularly effective at correcting failures on medium queries. This is likely because medium queries contain a high concentration of logic-related errors, which are well-suited to correction through enhanced reasoning. In contrast, easy queries tend to involve edge cases, while hard queries are often dominated by syntax issues and execution errors, where reasoning alone is insufficient to guarantee success. For example, OpenAI’s o4-mini corrects 25.36% more errors over GPT-4o in medium queries than hard or simple ones. However, the extent of improvement varies considerably across model families, highlighting that the effectiveness of reasoning depends not only on the presence of reasoning mechanisms but also on how well they are integrated and executed. We also conducted prompt ablation experiment and the results are provided in Appendix B.
5 Case Study
We observe a high frequency of logic errors across all models and test cases. Therefore, we focus our analysis on this category. Below are representative test cases that specifically highlight different forms of logical failure.
Figure 3: Example test case failures: over-reasoning (left) and sequential filtering (right).
  • Over-reasoning (Code and ReAct). Agents frequently demonstrate over-reasoning by performing unnecessary operations that go beyond the requirements of the prompt. This often stems from misinterpreting implicit cues. In direct code generation, agents tend to apply extra filters that overly constrain the result set, introducing conditions not specified in the query. In the ReAct setting, the model continues to invoke tools and apply filters even after all constraints have been met, lacking a clear signal for termination. As illustrated in Figure 3(left), the ReAct agent first retrieves parking capacity correctly, then unnecessarily proceeds to filter for the presence of shopping malls and many other unrelated tools. The agent accumulates irrelevant actions until it reaches the step limit and fails to return an answer, despite having the correct information early on.
  • Sequential Filtering (Code). In complex queries involving multiple chains of constraints, the model often applies filters sequentially rather than evaluating the logic holistically. As shown in Figure 3(right), the query contains an OR condition between competitor count and population. The correct approach (left) evaluates both branches independently before merging the results. However, the LLM agent (right) applies filters sequentially, first by competitor count and then by population, prematurely discarding zones that could have satisfied the second condition. This flawed execution leads to valid zones being excluded, revealing a logical failure where sequential filtering overrides the intended logical structure.
6 Related Work
LLM Reasoning. Reasoning represents a key capability of advanced LLMs, enabling them to decompose tasks into smaller subproblems, perform structured search, and make decisions for complex problem-solving. Various techniques have been proposed to enhance the reasoning abilities of LLMs. Prompting methods wei2022chain ; yao2023tree ; besta2024graph , such as chain-of-thought (CoT) prompting wei2022chain , encourage models to generate intermediate and structured reasoning steps before producing final answers. These sequential reasoning patterns can be further extended into more expressive structures, such as trees yao2023tree , for handling more complex reasoning workflows. Beyond prompting, recent research has shown that reasoning capabilities can also be distilled into LLMs’ generation itself through post-training strategies guo2025deepseek ; snell2024scaling ; brown2024large ; zelikman2024quiet . In particular, reinforcement learning is often used to fine-tune models to produce high-quality reasoning trajectories, such as CoT-style paths guo2025deepseek ; havrilla2024teaching ; kumar2024training ; carta2023grounding . In parallel, agentic approaches offer an alternative paradigm, equipping LLMs with tools, external memory, and planning mechanisms to solve complex tasks wiesinger2024agents ; yao2023react ; shinn2023reflexion ; zhao2024expel ; shen2023hugginggpt ; schick2023toolformer . For instance, the ReAct framework yao2023react enables in-context learning by prompting models to engage in iterative thought-action-observation cycles. In this work, we systematically evaluate both LLMs and reasoning-enhanced strategies on practical real-world reasoning tasks through our proposed benchmark.
Benchmarks for LLMs and Agents. Benchmarks play an essential role in the development and evaluation of LLMs. Fundamental language capabilities are typically assessed through general-purpose benchmarks targeting language understanding and factual question answering srivastava2022beyond ; wang2024mmlu ; wang2024mmlu ; rein2024gpqa . In addition, domain-specific benchmarks have been introduced to evaluate LLMs’ proficiency in specialized areas such as biology chen2023bioinfo and law fei2023lawbench . As reasoning has emerged as a central focus in recent LLM research, benchmarks tailored specifically to reasoning have become increasingly important. Given the complexity of evaluating LLM reasoning, existing benchmarks often focus on domains with well-defined verifiers, where checking the correctness of a solution is significantly easier than generating it. Two such domains–mathematics and programming–have become prominent battlegrounds for reasoning-oriented LLMs math500 ; quan2025codeelo ; aime2024 , exemplified by benchmarks such as MATH500 math500 and CodeForces quan2025codeelo . However, constructing benchmarks for practical reasoning tasks remains challenging, largely due to the difficulty of defining objective metrics and the high cost of manual annotation xie2024travelplanner . In contrast, our benchmark evaluates LLMs and agents on real-world site selection reasoning while maintaining full scalability through automatic verification, requiring no human annotation. This enables robust, reproducible, and extensible evaluation of reasoning capabilities in realistic decision-making contexts.
7 Conclusion
In this paper, we present LocationReasoner, a benchmark designed to evaluate LLMs and agents on real-world site selection reasoning tasks, supported by a sandbox environment enriched with built-in tools and curated datasets. Through systematic evaluation of both general-purpose and reasoning-oriented LLMs from multiple providers, along with various agentic strategies, we find that current models perform poorly–with even the most advanced reasoning LLM failing to solve over 30% of the queries. Our analysis identifies key bottlenecks, particularly in over-reasoning and the challenges of sequential problem-solving, which limit the effectiveness of LLMs in complex real-world reasoning scenarios. LocationReasoner serves as a scalable and extensible testbed: new constraints can be added and automatically validated without the need for human annotation. We hope LocationReasoner will drive progress in enhancing the reasoning capabilities of LLMs and contribute to the advancement of general intelligence for solving practical, high-stakes problems.
Appendix A Reasoning Architecture Comparison
Direct code generation consistently outperforms ReAct/Reflexion because of the fundamental differences in how each paradigm approaches logical reasoning. Below, we highlight two critical architectural dimensions: holistic vs. chunked planning and linear vs. nonlinear reasoning.
Holistic vs. Chunked Planning
Code generation promotes holistic planning. The model sees the entire query and designs a solution that integrates all constraints into one coherent logic block. It defers execution until the entire logical structure is assembled and performs filtering or aggregation after all conditions are accounted for. On the contrary, ReAct operates in chunks where each TAO cycle is focused on a single subproblem. The agent reasons incrementally, using only its most recent observations to decide what to do next. While this can help with short-term reactivity, the model loses sight of the broader goal. Instead of building a global solution from the outset, ReAct focuses on local decisions by evaluating one tool at a time without fully understanding how each TAO interacts, which results in fragmented logic, redundant actions, and plans that satisfy isolated constraints but fail as a cohesive whole.
Nonlinear vs. Linear Reasoning
Direct code generation follows a nonlinear reasoning approach. The model has the flexibility to write logic in any order, reuse variables fluidly, and jump between intermediate steps as needed. This flexibility allows the agent to integrate information from different sources without being tied to a rigid step-by-step flow. In contrast, ReAct enforces a linear reasoning pattern. The agent processes the query as a series of discrete TAO steps. Each step is tightly coupled to the previous one, which limits the model’s ability to reorganize its plan or make global optimizations. This structure makes ReAct more prone to local errors and brittle when reasoning requires backtracking or dynamic reorganization.
Appendix B Prompt ablation
To investigate whether prompt design can improve accuracy and reasoning ability, we conduct a prompt ablation study by adding explicit guidance to the instruction. We augment the prompt with phrases such as “think holistically,” “consider edge cases,” and “do not discard zones early.” This modification is evaluated on hard queries using the Gemini model family. The revised prompts are referred to as v2. As shown in Table 4, both Gemini 1.5 and Gemini 2.5 exhibit performance improvements with the revised prompting strategy. The perfect pass rate for Gemini 2.5 increases from 45.10% to 49.02%, a relative improvement of 8.7%, while Gemini 1.5 has a relative improvement of 4.9%. Similar trends are observed across precision, recall, and F1 score, with Gemini 2.5 showing especially notable gains. These findings suggest that better prompting can meaningfully enhance agent reasoning in complex queries where logical missteps and premature filtering are common failure modes. The pronounced improvement on Gemini 2.5 illustrates that stronger models are more responsive to strategic instruction tuning.
Table 4: Performance metrics on hard queries across Gemini model versions.
Model Delivery Rate Perfect Pass Precision Recall F1 Score
Gemini 2.5 83.33% 45.10% 0.3300 0.3500 0.3200
Gemini 2.5 v2 93.14% 49.02% 0.3792 0.3663 0.3437
Gemini 1.5 46.08% 19.61% 0.1300 0.0800 0.0900
Gemini 1.5 v2 43.14% 20.58% 0.1402 0.0901 0.0927
Appendix C Case Study: Dubai
Table 5: In-house tools for the Dubai dataset.
Category Tool Name Description
Loader Tool load_dubai_dataset Loads the Dubai dataset and returns a DataFrame named dubai_df.
Filter Tools dubai_near_the_sea Returns locations where the "Near the Sea" attribute is True.
dubai_community Returns locations where the "Community" attribute is True.
dubai_business Returns locations where the "Business" attribute is True.
dubai_tourist Returns locations where the "Tourist" attribute is True.
dubai_affordable Returns locations where the "affordable" attribute is True.
To evaluate the generalizability of our framework beyond the original setting, we apply the same task structure to a new urban context: the city of Dubai. The Dubai dataset consists of individual locations rather than zones, and all filtering functions return Boolean values. Specifically, the agent reasons over five simplified binary filters: near-to-sea, community, business, tourist, and affordable. Unlike the original in-house tools, which involved numerical thresholds, spatial computations, and logical operators over continuous data, the Dubai tools rely solely on simple DataFrame merges and categorical truth values, which is a less mathematically intensive reasoning environment.
We generated 30 natural language queries for this case study, evenly split across three difficulty levels that parallel those in the original benchmark. We evaluated both code generation models (Gemini 1.5 and Gemini 2.5) and agentic workflows powered by OpenAI’s GPT-4o under ReAct and ReAct + Reflexion. While the core task design remains unchanged, the shift in context and simplification of underlying functions introduces a new dimension for assessing model adaptability.
The simplified nature of the Dubai dataset led to a modest improvement in overall performance across models. In particular, the benchmark improved the performance of the Gemini 2.5 model, which achieved an overall F1 score of 0.88 and a perfect pass rate of 83 33%. This suggests that reduced complexity in tool use and data structure lowers the barrier for correct execution. However, despite these favorable conditions, the ReAct + Reflexion agent continued to underperform, indicating persistent challenges in planning capabilities.
ReAct + Reflexion achieved only a 53.33% delivery rate and an overall F1 score of 0.52, underperforming even the base ReAct agent. Its perfect pass rate (50.00%) matched that of the simpler ReAct agent. The additional self-reflection steps does not translate to consistency or execution quality. This persistent gap raises questions about the reliability of reflection-based mechanisms in environments where reasoning does not benefit from iterative correction or deep planning, but instead relies on simple filter chaining. These results suggest that while Reflexion may offer advantages in complex or abstract task settings, it may not be well-suited to binary, logic-oriented tasks such as those exemplified by the Dubai benchmark.
Table 6: Performance metrics on the Dubai dataset by model and difficulty level
Model Difficulty Delivery Rate Perfect Pass Precision Recall F1 Score
Code Generation
simple 100% 90% 1 0.92 0.93
medium 60% 50% 0.5 0.5 0.5
hard 70% 20% 0.2 0.2 0.2
Gemini 1.5 overall 76.67% 53.33% 0.57 0.54 0.54
simple 100% 90% 1 0.92 0.93
medium 90% 100% 1 1 1
hard 100% 60% 0.82 0.67 0.7
Gemini 2.5 overall 96.67% 83.33% 0.94 0.86 0.88
Agentic Workflow
ReAct (gpt-o4) sim 100.00% 50.00% 0.60 0.54 0.56
med 100.00% 40.00% 0.40 0.40 0.40
hard 80.00% 20.00% 0.20 0.20 0.20
overall 93.33% 36.67% 0.40 0.38 0.39
ReAct + Reflexion (gpt-o4) sim 100.00% 50.00% 0.63 0.53 0.55
med 40.00% 70.00% 0.70 0.70 0.70
hard 20.00% 30.00% 0.30 0.30 0.30
overall 53.33% 50.00% 0.54 0.51 0.52